In Anchorage, collisions of a vehicle with a moose are so commonthat they are referred to with the abbreviation MVC. Suppose a1050 kg car slides into a stationary520 kg moose on a very slippery road,with the moose being thrown through the windshield (a common MVCresult).
(a) What percent of the original kinetic energyis lost in the collision to other forms of energy?
%

(b) A similar danger occurs in Saudi Arabia because ofcamel-vehicle collisions (CVC). What percent of the originalkinetic energy is lost if the car hits a 270 kg camel?
%

(c)Generally, does the percent loss increase or decrease if the animalmass decreases?
increase
decrease

Is this an elastic collision or a completelyinelastic collision? Did you find the kinetic energy of the carbefore and the kinetic energy of the car + moose after? Did youconvert the speed to m/s?
Answer #1
Since momentum is conserved you can write:
   M vi =   (M +m)vf         where I used M for the car and m for the moose.
If you square both sides of the equation and multiply by(1/2), then you get
   (1/2) M2vi2    =   (1/2) (M+ m)2vf2             
Notice that this is almost the initial and final kineticenergies... but not quite. There is an extra factor of mass on eachside. Since   KE is (1/2) mv2    we can write each side as the KEtimes mass or
     M * Ki   = (M + m ) Kf
Then we can get the ratio of K final to K initial
    Kf / Ki = M / (M+m) = 1050 / (1050 + 520) = 0.6688  
this means 66.88 % of the KE is kept by the car andmoose... so 33.12% is "lost"
Similarly, for the car and camel
      Kf /Ki = M / (M+m) = 1050 / (1050 +270) = 0.7955  
this means 79.55 % of the KE is kept by the car andcamel... so 20.45% is "lost"